Some X, Y data are positively correlated with r = 0.5. Also:  

x̄ = 4 Sx = 2
ȳ = 6 Sy = 3

a. Find the equation for the regression line

b. Predict Y for X = 8

c. Find the coefficient of determination and interpret its meaning.
Answer #1

@ Slope : β., r Ï 3 Intercept ß - y - 3 X =6-0.75*4. 3 . The equastion ^or the vegression line . y= 3 +0.75 χ O For simple